OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. R(3, 3), S(-6, -6)

OpenStudy (anonymous):

first find slope of the line with this formula \[m=\frac{ y _{R} - y _{S} }{ x_{R}-x _{S} }\]

OpenStudy (anonymous):

@amirreza1870 I got 0...

OpenStudy (anonymous):

no.\[\frac{ -6-3 }{ -6-3 }=\frac{ -9 }{ -9 }=1\]

OpenStudy (anonymous):

oh ok and then what do I do?

OpenStudy (anonymous):

then use one of points for example R .and use this formula . \[y-y _{R}=m(x -x_{R})\]

OpenStudy (anonymous):

I got 3+6 = 1(3+6) is that correct? then I got 9= 1(9)

OpenStudy (anonymous):

no. only use R point y-3=1(x-3) y-3=x-3 y=x

OpenStudy (anonymous):

@amirreza1870 I'm not a genius in geometry :( can you help me

OpenStudy (anonymous):

@amirreza1870 is it x-y= 0?

OpenStudy (anonymous):

yes x-y=0 so x=y

OpenStudy (anonymous):

Omg! Thank you sooo muchhh